# Episodic Contextual Bandits with Knapsacks under Conversion Models

### v1

@@ A.2 Examples of Confidence Bound (CB) Oracles
While we primarily focus on contextual models on the conversion probability $\rho(\theta,a)$, in certain applications such as first price auction (see Section A.1.3) only the deterministic outcomes $r(\theta,a)$ and/or $d(\theta,a)$ depends on the context $\theta$, but not $\rho(\theta,a)$. In this case, the conversion probability $\rho(\theta,a)$ is modeled by a non-contextual $(A-1)$-armed bandit model.
@@ A.2.2 Contextual Generalized Linear Bandits [FilippiCGS10, RenZK22]
where $\tilde{O}(\cdot)$ hides a multiplicative factor of $\log^{3/2}(dTH/\kappa_{f}\delta)$.
@@ A.2.3 Refined Regret Bounds for Contextual Logistic Bandits [FauryACF2020, LiS22]
Given the above $\hat{\mu}$, we plug in the value and derive $\hat{P}(\theta,a)=f(\phi(\theta_{n},A_{n})^{\top}\hat{\mu})$ as the estimated conversion probability. Based on the calculation in Lemma 1 of [LiS22], and the fact that $\|B(1)\|_{2}=1$, we define
@@
where $\tilde{O}(\cdot)$ hides a multiplicative factor of $\log^{3/2}(dTH/\kappa_{f}\delta)$.
@@ A.2.4 Contextual Kernelized Bandits [ValkoKMFC13, ChowdhuryG17]
Following [ValkoKMFC13, ChowdhuryG17] and being motivated by Gaussian processes, we assume that $p\in\Xi$, where $\Xi$ consists of functions that have a small RKHS norm. We provide a more precise definition of $\Xi$ in the following discussion. We let $\Xi$ be an RKHS of functions that map $\Phi\rightarrow[0,1]$. The RKHS $\Xi$ is associated with a positive semi-definite kernel function $k:\Phi\times\Phi\rightarrow\mathbb{R}$. The kernel $k$ induces an inner product operator $<\cdot,\cdot>_{k}$ that obeys the reproducing property: $\bar{p}(\phi)=<\bar{p},k(\phi,\cdot)>_{k}$ for all $\bar{p}\in\Xi$ and all $\phi\in\Phi$, and that $k(\phi,\phi^{\prime})\leq 1$ for all $\phi,\phi^{\prime}\in\Phi$. Common examples of such kernels are the Squared Exponential kernel and the Matern kernel, respectively defined as
@@ Appendix B Proofs
Before we start the proof, we recall the event ${\cal F}_{t,\delta^{\prime}}$ defined in Lemma 3.4, and the event ${\cal G}_{t,\delta_{t}}$ defined in (6). The event ~$\tilde{\cal E}_{h,t,\delta^{\prime}}$ is defined in (54) in this Appendix section. In addition, $(\star_{t})=(\diamondsuit_{t})+(\dagger_{t})+(\ddagger_{t})$. The quantities $(\diamondsuit_{t}),(\dagger_{t}),(\ddagger_{t})$ are defined in the forthcoming (46, 56, 60) respectively, and the inequality $\Pr(\sum^{T}_{t=1}(\star_{t})\leq 2r_{\text{max}}(2L+1)\sqrt{HT\log(2/\delta^{\prime\prime})})\geq 1-3\delta^{\prime\prime}$ for any $\delta^{\prime\prime}\in(0,1)$ is justified by a union bound on the forthcoming (49, 58,62).
@@
where the event ~$\tilde{\cal E}_{h,t,\delta^{\prime}}$ is defined in the forthcoming (54), and the error terms $(\dagger_{h,t},\ddagger_{h,t})$ are defined in the forthcoming (56, 60) respectively. To prove (50), consider a fixed time step $(h,t)$ where $h\in[H]$:
@@
~$\tilde{\cal E}_{h,t,\delta^{\prime}}=\left\{\frac{1}{|{\cal J}_{t}|+M}\sum_{s\in{\cal J}_{t}\cup\{-1,\ldots,-M\}}\hat{V}_{h,t}(b,\theta_{h,s})-\mathbb{E}_{\theta_{h}\sim\Lambda_{h}}[\hat{V}_{h,t}(b,\theta_{h})~{}|~{}\hat{V}_{h,t}]\right.$
@@ Appendix C Numeric Experiment
At episode $t$, round $h-1$, the Oracle for algorithms SquareCBwK and Logistics-UCB1, would use the history $H=\{\theta_{h^{\prime},t^{\prime}},A_{h^{\prime},t^{\prime}}\}_{h^{\prime}=1,t^{\prime}=1}^{h-1,t-1}\cup\{\theta_{h^{\prime},t},A_{h^{\prime},t}\}_{h^{\prime}=1}^{h-1}$ to train the model and output the predicted conversion probability or predicted mean reward. Specifically, we implement Logistics-UCB1 by taking $B_{T}=B$, to enhance its numeric performance.

### v2

@@ A.2 Examples of Confidence Bound (CB) Oracles
While we primarily focus on contextual models on the conversion probability $\rho(\theta,a)$, in certain applications such as first price auction (see Section LABEL:sec:auction) only the deterministic outcomes $r(\theta,a)$ and/or $d(\theta,a)$ depends on the context $\theta$, but not $\rho(\theta,a)$. In this case, the conversion probability $\rho(\theta,a)$ is modeled by a non-contextual (A1)(A-1)-armed bandit model.
@@ A.2.2 Contextual Generalized Linear Bandits [FilippiCGS10, RenZK22]
[…]
@@ A.2.3 Refined Regret Bounds for Contextual Logistic Bandits [FauryACF2020, LiS22]
[…]
@@
[…]
@@ A.2.4 Contextual Kernelized Bandits [ValkoKMFC13, ChowdhuryG17]
Following [ValkoKMFC13, ChowdhuryG17] and being motivated by Gaussian processes, we assume that $p\in\Xi$, where $\Xi$ consists of functions that have a small RKHS norm. We provide a more precise definition of $\Xi$ in the following discussion. We let $\Xi$ be an RKHS of functions that map $\Phi\rightarrow[0,1]$. The RKHS $\Xi$ is associated with a positive semi-definite kernel function $k:\Phi\times\Phi\rightarrow\mathbb{R}$. The kernel $k$ induces an inner product operator $\cdot,\cdot>_{k}$ that obeys the reproducing property: $\bar{p}(\phi)=<\bar{p},k(\phi,\cdot)>_{k}$ for all $\bar{p}\in\Xi$ and all $\phi\in\Phi$, and that $k(\phi,\phi^{\prime})\leq 1$ for all $\phi,\phi^{\prime}\in\Phi$. Common examples of such kernels are the Squared Exponential kernel and the Matern kernel, respectively defined as
@@ Appendix B Proofs
Before we start the proof, we recall the event ${\cal F}_{t,\delta^{\prime}}$ defined in Lemma 3.4, and the event ${\cal G}_{t,\delta_{t}}$ defined in (6). The event $\tilde{\cal E}_{h,t,\delta^{\prime}}$ is defined in (54) in this Appendix section. In addition, $(\star_{t})=(\diamondsuit_{t})+(\dagger_{t})+(\ddagger_{t})$. The quantities $(\diamondsuit_{t}),(\dagger_{t}),(\ddagger_{t})$ are defined in the forthcoming (46, 56, 60) respectively, and the inequality $\Pr(\sum^{T}_{t=1}(\star_{t})\leq 2r_{\text{max}}(2L+1)\sqrt{HT\log(2/\delta^{\prime\prime})})\geq 1-3\delta^{\prime\prime}$ for any $\delta^{\prime\prime}\in(0,1)$ is justified by a union bound on the forthcoming (49, 58,62).
@@
where the event $\tilde{\cal E}_{h,t,\delta^{\prime}}$ is defined in the forthcoming (54), and the error terms $(\dagger_{h,t},\ddagger_{h,t})$ are defined in the forthcoming (56, 60) respectively. To prove (50), consider a fixed time step $(h,t)$ where $h\in[H]$:
@@
$\tilde{\cal E}_{h,t,\delta^{\prime}}=\left\{\frac{1}{|{\cal J}_{t}|+M}\sum_{s\in{\cal J}_{t}\cup\{-1,\ldots,-M\}}\hat{V}_{h,t}(b,\theta_{h,s})-\mathbb{E}_{\theta_{h}\sim\Lambda_{h}}[\hat{V}_{h,t}(b,\theta_{h})~|~\hat{V}_{h,t}]\right$.
@@ Appendix C Numeric Experiment
At episode $t$, round h1h-1, the Oracle for algorithms SquareCBwK and Logistics-UCB1, would use the history $H=\{\theta_{h^{\prime},t^{\prime}},A_{h^{\prime},t^{\prime}}\}_{h^{\prime}=1,t^{\prime}=1}^{h-1,t-1}\cup\{\theta_{h^{\prime},t},A_{h^{\prime},t}\}_{h^{\prime}=1}^{h-1}$ to train the model and output the predicted conversion probability or predicted mean reward. Specifically, we implement Logistics-UCB1 by taking $B_{T}=B$, to enhance its numeric performance.